# Detecting charge transfer at defects in 2D materials with electron ptychography

### v4

@@ Abstract
Charge transfer between atoms is fundamental to chemical bonding but has remained very challenging to detect directly in real space. Atomic-resolution imaging of charge density is not sufficient by itself, as the change in the density due to bonding is very subtle compared to the total local charge density. Sufficiently high sensitivity, precision and accuracy are required, which we demonstrate here for the detection of charge transfer at defects in two-dimensional WS2 via high-speed electron ptychography and its ability to correct errors due to residual lens aberrations.
@@
Phase-contrast imaging methods, on the other hand, are known for being highly dose efficient, and indeed, local atomic-scale detection of charge redistribution due to bonding has been demonstrated in high-resolution transmission electron microscopy (HRTEM) for both single-layer hexagonal boron nitride and doped graphene [10]. However, revealing the charge redistribution in HRTEM required careful use of specific contrast transfer function (CTF) conditions with large defocus values, significantly degrading image resolution. Furthermore, due to the oscillatory CTFs of HRTEM, it is often difficult to interpret even for structural imaging, let alone disentangling charge transfer from that of the atomic structure. Fortunately, STEM phase-contrast imaging has improved greatly in the past decade. Multiple phase-imaging modes can be performed without the need for aberrations and in parallel with -contrast ADF, providing easily interpretable images of structures at the maximum resolution of the microscope. Atomic-resolution charge-density imaging has been performed with differential phase contrast (DPC), scattering center of mass (CoM) [11] and off-axis holography [12], but these techniques have not been sufficient to detect charge transfer due to bonding.
@@
Here we employ high-speed [13] focused-probe single-sideband (SSB) electron ptychography with post-collection aberration correction [14] to provide the necessary precision and accuracy to detect charge transfer directly in WS2, including at individual point defects. In addition to the high-speed and dose-efficiency of our method allowing us to outrun damage, the ptychographic aberration correction overcomes the problem of residual aberrations that has so far prevented CoM-based methods from achieving valence imaging [15]. Accounting for residual aberrations, as we do here, is essential because they can completely obscure the effect of charge transfer, and if ignored may result in an entirely incorrect interpretation. Also crucial is the use of our new parameter-based quantification method [16] that explicitly accounts for the CTF of the imaging modality and overcomes artefacts due to the sample configuration such as mistilt.
@@
The problem of residual aberrations is illustrated in Fig. 1 with experimental data of WS2 from a modern aberration-corrected STEM (see Methods in Supplemental Material). Aberrations are particularly problematic when quantifying phase images [17]. Two SSB images from the same dataset are shown, one with ptychographic aberration correction and one without. Although both were acquired using the electron-optical aberration corrector, the lattice polarity in the SSB image without the additional ptychographic post-collection correction is reversed because of the residual aberrations. With such a large effect on the phase from relatively small aberrations, interpreting phase images from either ptychography or CoM-based methods, including DPC, is fraught with danger, and reliably doing so in terms of the subtle effects of charge transfer is essentially untenable.
@@
For very thin specimens, such as 2D materials, without aberrations the phase should be flat in the double-disk overlaps in probe reciprocal space, a pair of which is shown for each SSB image in Fig. 1. This is clearly not the case without ptychographic aberration correction. We use singular value decomposition (SVD) to identify the aberrations present and counteract the phases of these aberrations, resulting in a nearly flat phase in the double-overlap regions, confirming the removal of the residual aberrations [14].
@@
To understand the role of charge transfer in the observed contrast, we employ first-principles DFT calculations and SSB ptychography image simulations based on both DFT and independent atom model (IAM) potentials. 4D-STEM datasets were simulated using the multislice algorithm with the abTEM code [18], with probe parameters and sampling set to match our experimental conditions (see Methods). The results show that the contrast between the sublattice sites depends on the convergence angle of the incoming probe. For the 20 mrad convergence angle used in Fig. 1, the IAM-based S2 phases are higher than those of the W sites. However, once bonding is included by using a scattering potential simulated with DFT, the contrast reverses, with the W sites having higher phase than the S2 sites. This is exactly what is seen in Fig. 1c once post-collection aberration correction is applied to the experimental ptychographic image.
@@
To quantify the influence of charge redistribution from the ptychographic images, we need to assign a phase to each of the atomic sites. Various approaches have been established for quantifying the atomic intensities of ADF images, including local maxima [7], the integration of Voronoi cells [19], Gaussian fits [20, 21], as well as template matching [22]. However, these methods do not work well for ptychography since its CTF induces a stronger nonlinear dependence on the local atomic environment [16]: neighbouring atoms can actually decrease the phase at a lattice site, which is very different to the additive behavior of probe tails in ADF imaging. This introduces a strong dependence not only on the neighboring atoms but also on sample tilt, which must be accounted for to correctly quantify such phase images. Notably, this behavior also occurs when CoM- or DPC-based images are high-pass filtered, as they often are.
@@
We recently developed a quantification method that accounts for the contrast transfer function of these methods and is robust to both sample tilt and source size [16]. This is important here because mistilt is particularly difficult to eliminate in 2D materials and can cause significant errors in quantification, in addition to residual aberrations. With this method, the quantified phase of an atomic column is independent of both the proximity and type of the neighboring columns.
Detecting charge transfer requires that the precision of the phase measurement is better than the difference caused by the transfer, which is far more challenging than simply mapping the charge density. However, due to its effect on screening the nucleus, charge transfer can result in an order of magnitude larger change in the phase than in the potential itself (SFig. 2,3), substantially aiding its detection. Experimentally, sufficient precision is either achieved with high doses, combining measurements from multiple identical atomic sites, or both. As the dose is limited by the specimen stability, averaging over multiple sites is often unavoidable. Based on simulated estimates of the dose-dependent precision of the phase extraction of charge transfer due to bonding for WS2 (SFig. 5), we estimate doses of eÅ and above to be sufficient when averaging over three pairs of sites.
Fig. 2 compares experimental and simulated SSB ptychography of pristine WS2 at a dose of eÅ. Line profiles taken across the W and S2 sites are shown in Fig. 2b. The experimental line trace is an average of 15 pairs of sites, which at this dose should produce a very high statistical significance. Including bonding via DFT clearly alters the relative phases of the two sublattices in comparison to the IAM simulation. The DFT potential is decreased at the S2 sites compared to the IAM because of the additional screening from the transfer of the 1.36 electrons from the tungsten atoms to the neighboring sulfur atoms; this increase in negative charge increases the screening of the positively charged nuclei at the sulfur sites, resulting in a lower overall positive potential and a reduced phase in ptychographic imaging. As can be seen, the experimental line trace matches the DFT simulation with bonding very well indeed, and is not compatible with the neglect of bonding in the IAM result.
@@
Fig. 2d shows that the ratio of the phases of the sites that in Fig. 2c are quantified for 20 mrad depends on the probe convergence angle. This is because the convergence angle determines the strength at which different spatial frequencies are transferred in the SSB method [23, 24], which in turn determines the relative strengths at which the features of the charge density distribution appear in the phase images. However, the fact that we can see these clear differences demonstrates that the frequencies transferred by these conditions do contain information on the charge transfer induced by bonding. We emphasize that such conditions are all in the range typical for high-resolution STEM imaging. This is in contrast to conventional HRTEM, where the imaging conditions relevant to detecting charge transfer are generally not those desired for imaging the structure itself. Including atomic vibrations using thermal diffuse scattering (TDS, see Methods) slightly decreases the S2 to W ratio compared to neglecting it, and the experimental uncertainty, estimated from the statistical site-by-site phase variation dominated by the limited signal, overlaps both values with bonding included.
@@
The simplest defect in transition metal dichalcogenides is a chalcogen vacancy. For WS2 this can be realized by removing a single S atom, leaving both a S vacancy and one remaining S atom at the former S2 site (V). These vacancies produce a significantly reduced phase in SSB reconstructions, as shown in the DFT-based SSB images in Fig. 3a, where three S vacancy sites are distributed around the central W atom (V). This data also showcases the ability of ptychography to reveal light atoms next to heavy elements [14, 25]: the S sites are in general not readily visible in the high-angle ADF signal under these conditions, and thus S vacancies are only discernible in the phase images (see also SFig. 6,7).
@@
Interestingly, we observe a higher W site phase close to the S vacancies, with the intensity of the W depending on the density of the vacancies in the area. The reason for this is two-fold: firstly, the CTF of the SSB reconstruction results in a negative halo that reduces the phase of atoms close to other atomic sites. Importantly, our phase optimization method takes this into account, allowing us to detect the second effect, the charge transfer.
@@
To systematically study the charge transfer at defects, we simulated a region of WS2 with an increasing numbers of S vacancies using both IAM and DFT potentials. Fig. 3a shows the DFT-based SSB phase image with three monovacancies (V). The difference between IAM- and DFT-based potentials, shown in Fig. 3b, increases at the W sites next to vacancies and further increases with the density of defects (see also SFig. 8).
@@
This increase would be surprising if one assumes that defects reduce charge transfer due to W atoms having fewer S neighbors. If that were the case, one would see a smaller phase shift in the DFT-based simulations, which is not what is observed. Fig. 4 shows variation of the phase ratio of the central W site to a neighboring S vacancy site () with respect to the density of vacancies. In the extreme three-divacancy case (V), the W atom has lost all its nearest neighbors and would therefore seem more similar to an independent atom, for which one might expect a negligible difference between DFT and IAM. However, this is not the case (see simulated SSB images and projected all-electron potentials in SFigs. 7,8) and this highly defected case is in fact further from the IAM result with a higher ratio than the other results using the DFT potentials. The IAM results show a negligible variation between the configurations as they should. The fact that S vacancies produce occupied states at the defect sites [26, 27] does not explain the observed contrast, and additional studies have shown charge redistribution occurs towards the defect sites [28].
@@
To understand this phenomenon, we plot the three-dimensional (pseudo-)electron valence densities of each configuration in SFig. 9, where a different picture emerges. The line profiles over the W site close to the defects indeed show a lower density of valence electrons when increasing the number of defects, in agreement with the increase of phase we observe at this site. The electrons are redistributed towards the next W sites, which creates W-W bonding that is stronger the more S vacancies are involved in the system. Thus we need to consider longer-range charge rearrangement than merely the first nearest neighbors to understand the full picture.
@@
The relative difference between IAM and DFT can reach up to 10% (cf. Fig. 4c), which is higher than in the pristine case and one might thus conclude that detecting charge transfer is easier at defects. However, defect configurations are often unique and complex and therefore a statistical assessment such as analysing multiple sites is very challenging with each atom potentially undergoing different amounts of charge transfer. Secondly, the dose one can use to probe defective configurations is usually lower, due to their greater propensity to damage, increasing the uncertainty per site. In this case the dose we could use before damage was only eÅ.
@@
To ascertain if we can detect charge transfer at defects experimentally, we used electron irradiation to obtain a variety of vacancy densities and atomic configurations in the sample. A representative area is shown in Fig. 3d (full view in SFig. 5). Fig. 3c shows the extracted phases of the W atoms in the pristine area (blue) and the W atoms close to vacancies, labeled as (dark purple). Indeed, the calculated mean value for shows a small but detectable shift in the DFT simulations compared to pristine W atoms. Both types of simulations were treated with Poisson noise in the 4D diffraction data corresponding to the experimental dose. Note that the shift is very small as the simulated structure only contains monovacancies (cf. SFig. 6a,b). The shift would be significantly higher if this specific area of interest also contained divacancies. This qualitative assessment also agrees with the experimental data, where the shift between and W is observable in Fig. 3e.
To quantify the effect, we calculated phase ratios between the atomic sites where the charge transfer occurs. We compared the ratio of the W site to vacancy site phases, and indeed find a significant range of ratios ( from 1.99 to 2.23) when analysing different areas in the sample. Fig. 4a and b show areas of the sample with different vacancy configurations and densities. The trend of charge transfer increasing with the amount of defects is also reflected in the experimental ratios (see Fig. 4c): Configurations 0 and 1 do not contain any divacancies and their ratios are in the lower range. Configurations 2–5 all contain divacancies and their ratios are clearly higher indicating a higher amount of charge transfer. Interestingly, configurations 2 and 4 are the only ones which contain neighboring divacancies, resulting in an even higher ratio (and charge transfer), again in agreement with the simulation analysis.

### v5

@@ Abstract
Electronic charge transfer at the atomic scale can reveal fundamental information about chemical bonding, but is far more challenging to directly image than the atomic structure. The charge density is dominated by the atomic nuclei, with bonding causing only a small perturbation. Thus detecting any change due to bonding requires a higher level of sensitivity than imaging structure and the overall charge density. Here we achieve the sensitivity required to detect charge transfer in both pristine and defected monolayer WS2 using the high dose efficiency of electron ptychography and its ability to correct for lens aberrations. Excellent agreement is achieved with first-principles image simulations including where thermal diffuse scattering is explicitly modeled via finite-temperature molecular dynamics based on density functional theory. The focused-probe ptychography configuration we use also provides the important ability to concurrently collect the annular dark-field signal, which can be unambiguously interpreted in terms of the atomic structure and chemical identity of the atoms, independently of the charge transfer. Our results demonstrate both the power of ptychographic reconstructions and the importance of quantitatively accurate simulations to aid their interpretation.
1 Introduction
@@
Phase-contrast imaging methods, on the other hand, are known for being highly dose efficient, and indeed, local atomic-scale detection of charge redistribution due to bonding has been demonstrated in high-resolution transmission electron microscopy (HRTEM) for both single-layer hexagonal boron nitride and doped graphene [10]. However, revealing the charge redistribution in HRTEM required careful use of specific contrast transfer function (CTF) conditions with large defocus values, significantly degrading image resolution. Furthermore, due to the oscillatory CTFs of HRTEM, it is often difficult to interpret even for structural imaging, let alone disentangling charge transfer from that of the atomic structure. Fortunately, STEM phase-contrast imaging has improved greatly in the past decade. Multiple phase-imaging modes can be performed without the need for aberrations and in parallel with -contrast ADF, which provides easily interpretable images of structures at the maximum resolution of the microscope. Atomic-resolution charge-density imaging has been performed with differential phase contrast (DPC), scattering center of mass (CoM) [11] and off-axis holography [12], but these techniques have not been sufficient to detect charge transfer due to bonding.
@@
Here we use high-speed [13] focused-probe single-sideband (SSB) electron ptychography with post-collection residual aberration correction [14] to provide the necessary precision and accuracy to detect charge transfer directly in WS2, including at individual point defects. In addition to the high speed and dose-efficiency of our method allowing us to outrun damage, the ptychographic aberration correction overcomes the problem of residual aberrations that has so far prevented CoM-based methods from achieving valence imaging at atomic resolution [15, 16]. Accounting for residual aberrations, as we do here, is essential because they can completely obscure the effect of charge transfer, and if ignored may result in an entirely incorrect interpretation. Also crucial is the use of our recent parameter-based quantification method [17] that explicitly accounts for the CTF of the imaging modality and overcomes artefacts due to the sample configuration such as mistilt.
Using the fact that the phase shift is directly proportional to the electric potential for thin samples, analogously to CoM imaging [18], we can take the Laplacian of the SSB phase image to obtain an image that is proportional to the charge density. This allows one to obtain an aberration-corrected charge density map, which is not possible in present CoM or DPC based methods. Here, we show that both ptychographic phase and charge density imaging can reveal the charge transfer, but analysing the phase images is more robust to noise than the charge density map itself because of its reliance on noise-amplifying differentiation [19].
2 Results and discussion
2.1 Pristine WS2
@@
The problem of residual aberrations is illustrated in Fig. 1 with experimental data from WS2 using a modern aberration-corrected STEM (see Methods). Aberrations are particularly problematic when quantifying phase images since even small values can alter the phase [20]. Two SSB images from the same dataset are shown in Fig. 1a, one with ptychographic residual aberration correction and one without. Although the data were acquired while using a carefully tuned 5th-order electron-optical aberration corrector, the lattice polarity in the SSB image without the additional ptychographic post-collection correction is reversed because of the residual aberrations, as illustrated by the line profiles in Fig. 1b. With such large effects on the phase images resulting from such relatively small aberrations, reliably interpreting uncorrected phase images from either ptychography or CoM-based methods, including DPC, to uncover the subtle effects of charge transfer is in our view essentially untenable.
@@
For very thin specimens such as 2D materials, the phase without aberrations should be flat in the double-disk overlaps in probe reciprocal space. The double-disk overlaps from two different spatial frequencies contained in the data are shown in Fig. 1c. Without ptychographic aberration correction, the double-disk overlaps of both spatial frequencies are not flat. We use singular value decomposition (SVD) to identify the aberrations present and counteract the phases of these aberrations, resulting in a nearly flat phase in each of the overlap regions, confirming the removal of the residual aberrations [14]. See Supplementary Material for additional information.
Fig. 1a also shows charge density maps calculated from the CoM and from the ptychographic phase images using differentiation. The raw CoM-based charge density map, the differentiated CoM (dCoM), is very noisy indeed and a Gaussian filter is required to make the structure apparent over the severe noise (see Supplementary Figure 2). The resulting filtered dCoM image shown in Fig. 1a is fairly similar to the SSB-based charge-density map without ptychographic aberration correction. The SSB-based charge-density map with the additional ptychographic aberration correction, however, shows a much higher-quality image. As we will show, our post-collection aberration-corrected ptychographic phase images and charge density maps show an excellent match with simulations that include charge transfer.
@@
To understand the role of charge transfer in the observed contrast, we simulated images with and without the effects of bonding. To do so we used potentials based on first-principles DFT and the independent atom model (IAM) as input to multislice 4D-STEM simulations using the abTEM code [21], with probe parameters and sampling set to match our experimental conditions (see Methods). Fig. 2 shows the ptychographic phase and charge density images simulated based on the DFT and IAM potentials. The figure also shows the differences between the DFT and IAM based images, i.e. the quantitative difference induced by the bonding. Compared to the IAM the DFT phase image has a higher phase on the W sites and a reduced phase on the S2 sites. Compared to the IAM based ptychographic charge density image, the DFT based result shows a higher charge density at the W sites and a reduced charge density at the S2 sites.
As the charge redistribution due to bonding contained in the DFT potential is the only significant difference between the DFT and IAM images, the differences between them are attributed to the charge transfer from W to S that occurs with electron redistribution due to bonding (see also SFigs. 3 and 4). Importantly, the magnitude of the difference due to bonding is around 10% of the maximum values at the atomic sites in the DFT and IAM images, enough to be detected at moderate electron doses (see Supplementary Material).
Fig. 2a also shows the result of subtracting the IAM phase and charge density images from the experimental images (scaling the theoretical ones to match the mean and standard deviation of the experimental images). The subtraction is very difficult and prone to error, especially because small mismatches in the alignment of the images will result in large values in the difference images. Despite the difficulty, a qualitative match with the theoretical image is obtained. The experimental difference image does however exhibit a magnitude three times higher than the theoretical difference, especially closer to the center of the hexagons. We expect this difference is mostly due to misalignment caused by experimental scan distortions, sample drift and mistilt, however residual aberrations which are not perfectly accounted for even with post-collection aberration correction may also contribute.
The phase differences between the IAM, DFT and experimental results are also reflected in the line profiles taken across the W-S2 dumbbells shown in Fig. 2c. An average of 45 dumbbells was used for the experimental profiles to provide high statistical significance. Interestingly, the charge density image provides less statistical significance than the phase image, as the difference image is noisier. This is likely because calculating the charge density map involves double differentiation, and differentiation is often associated with noise amplification. We therefore propose that analysing the phase is more reliable when it comes to charge transfer detection.
@@
To quantify the influence of charge redistribution from the ptychographic images, we need to assign a phase to each of the atomic sites. Various approaches have been established for quantifying the atomic intensities of ADF images, including local maxima [7], the integration of Voronoi cells [22], Gaussian fits [23, 24], as well as template matching [25]. However, these quantification methods do not work well for methods with low transfer of low spatial frequencies such as filtered CoM based imaging and ptychography, as they induce a nonlinear dependence on the local atomic environment [17]: the proximity of neighbouring atoms can actually decrease the intensity at which they appear in images, which is very different to the additive behavior of probe tails in ADF imaging. This introduces a strong image dependence not only on the atomic arrangement but also on sample tilt, which must be accounted for to correctly quantify such phase images, but is not accounted for in any of the usual image quantification methods designed for ADF imaging.
@@
To overcome this problem, we recently developed a quantification method that specifically accounts for the contrast transfer function (CTF) of the imaging method and is robust to both sample tilt and source size [17]. A model fitting the experimental data is initiated and a fast simulation using the convolution between the point-potential of the model and a kernel representing the CTF is optimized with respect to the experimental image. This is especially important here because mistilt is particularly difficult to eliminate in 2D materials and would cause significant errors in quantification, in addition to any uncorrected residual aberrations. With this method, the quantified phase of an atomic column is independent of both the proximity and type of the neighboring atoms. In addition to the SSB used presently, the quantification method can also be used for iterative ptychography and filtered CoM based imaging, which can also exhibit a negative halo around isolated atoms, as we demonstrated in reference [17].
Fig. 2d shows a histogram of the quantified phases of the W and S atoms in the experimental image. The phases of the two types of atoms are well separated. The ratio of the mean quantified phase values of S2 to W depends on the convergence angle as shown in Fig. 2e and, crucially, matches the results of the DFT simulations and not the IAM results.
@@
The variation with the convergence angle occurs because it determines the strength at which different spatial frequencies are transferred [26, 27]. This in turn determines the relative strengths at which the features of the charge density distribution appear in the phase images. The excellent agreement of the experiment at both convergence angles with the DFT results, and lack thereof for the IAM results, demonstrates that the frequencies transferred by these conditions contain the information on the charge transfer induced by bonding. We emphasize that such conditions are typical for high-resolution STEM imaging. This is in contrast to the situation in conventional HRTEM, where the imaging conditions relevant to detecting charge transfer are generally not those desired for imaging the structure itself. Including atomic vibrations using thermal diffuse scattering (TDS, see Methods) slightly decreases the S2 to W ratio compared to neglecting it, and the experimental uncertainty, estimated from the statistical site-by-site phase variation, dominated by the limited signal, overlaps both values with bonding included.
2.2 Defective WS2
@@
The simplest defect in transition metal dichalcogenides is a chalcogen vacancy. For WS2 such a monovacancy (VS) can be realized by removing a single S atom, leaving both a S vacancy and one remaining S atom at the former S2 site. Such vacancies produce a significantly reduced phase at the vacancy sites in SSB reconstructions, as shown in the DFT-based SSB image in Fig. 3a, where three S monovacancies (V3S) are distributed around the central W atom. Note that the S sites are in general not readily visible in the high-angle ADF signal under these conditions, and thus S vacancies are only discernible in the phase images (see also SFig. 7 and 8), highlighting the ability of ptychography to reveal light atoms next to heavy elements [14, 28].
@@
Interestingly, we observe a higher W site phase close to the S vacancies in our SSB images, with the intensity of the W depending on the density of the vacancies in the area. The reason for this is two-fold: the CTF, and the charge transfer itself. As we discussed earlier, the SSB CTF means that the phase of atoms is reduced in SSB images when they are brought near to each other. Therefore, removing an atom to form a vacancy significantly increases the phase of the neighboring sites in SSB images even for neutral atoms. It is therefore vital to take this CTF effect into account, as we do with our method of phase quantification, allowing us to detect the second effect, the charge transfer.
@@
To systematically study the charge transfer at defects, we simulated images of a region of WS2 with increasing numbers of S vacancies using both IAM and DFT potentials (SFig. 9). For the the V3S configuration shown in Fig. 3a, the difference between the IAM- and DFT-based image simulations is displayed in Fig. 3b. Keeping in mind the differences seen in pristine WS2, the effect of the vacancies on the difference is greatest at the W atom surrounded by the vacancies. Furthermore, we find that the phase of a W atom in WS2 increases in the DFT-based images relative to that of the IAM-based images proportionally with the density of S vacancies around the W atom. This is clearly seen in SFig. 9, with the DFT phase at the W atom increasing significantly more with increasing vacancy density than the IAM.
The variation of the quantified phase ratio of the central W site to a S vacancy site () with respect to the density of vacancies for the DFT- and IAM-based simulations is shown in Fig. 4c. Each subsequent configuration in the simulations has one less S atom. The ratios for the IAM simulations results are essentially flat, while the DFT ratios increase linearly with the density of vacancies. The flatness of the IAM result is exactly what is expected given our quantification method accounts for changes that are purely structural. As the main difference between the IAM and DFT images is the charge transfer in the DFT-based images, the DFT results show that the amount of charge transfer at a W site increases with the amount vacancies surrounding it and that this results in a higher quantified phase of the site.
@@
This increase in the charge transfer with vacancy density would be surprising if one assumes that vacancies reduce charge transfer due to the W atoms having fewer S neighbors. If that were the case, one would see a reduced phase in the DFT images relative to the the IAM images as the vacancy density is increased, rather than the observed increase. In the extreme three-divacancy configuration (V), the W atom has lost all its nearest neighbors and one might think is therefore more similar to an independent atom, and expect a smaller difference between the DFT and IAM results. However, this is clearly not the case, as seen in the simulated SSB images and projected all-electron potentials shown in SFigs. 9 and 10 and this highly defected case is in fact the furthest from the IAM result with a higher ratio than all the other results using the DFT potentials. The fact that S vacancies produce occupied states at the defect sites [29, 30] does not explain the observed contrast, and additional studies have shown charge redistribution occurs towards the defect sites [31].
@@
To understand this physical phenomenon, we plot the three-dimensional (pseudo-)electron valence densities of each configuration in SFig. 11, from which a different picture emerges. The line profiles from W sites adjacent to the defects show a lower density of valence electrons when increasing the number of vacancies around them, in agreement with the increase of phase we observe at this site. The electrons are redistributed towards the next W sites, which creates W-W bonding that is stronger the more S vacancies are involved in the system. Thus we need to consider longer-range charge rearrangement than merely the first-nearest neighbors to understand the full picture.
@@
The relative difference between IAM and DFT phases can reach up to 10% with vacancies (cf. Fig. 4c). This is a larger difference than in the pristine case, and one might thus conclude that detecting charge transfer is easier at defects. However, defect configurations are often unique and complex and therefore a statistical assessment analysing multiple sites is very challenging with each atom potentially undergoing different amounts of charge transfer. Further, the dose one can use to probe defective configurations is usually lower due to their greater propensity to damage, increasing the uncertainty per site. In this case the dose we could use before damage was only eÅ2.
@@
To ascertain whether we can detect charge transfer at defects experimentally, we used electron irradiation to obtain a variety of vacancy densities and atomic configurations in the sample. A representative area is shown in Fig. 3d, with the full field of view shown in SFig. 7. Fig. 3c shows histograms of the extracted phases of the W atoms in the pristine area (blue) and the W atoms adjacent to vacancies, labeled as (dark purple) for the IAM- and DFT-based simulations. Both types of simulations were treated with Poisson noise in the 4D diffraction data corresponding to the experimental dose. The calculated mean value for shows a significant increase in phase compared to pristine W atoms in the DFT simulations. Note that the structure analyzed here only contains monovacancies (cf. SFig. 8a,b); and that the phase shift due to charge transfer would be significantly higher if this specific area of interest also contained divacancies. This agrees with the experimental data, where an increase in the mean phase of the compared to that of the pristine W sites is indeed observable in Fig. 3e.
To further study the effect of the vacancies on the charge transfer experimentally, we calculated the phase ratio of the sites between which the majority of the charge transfer occurs, , using the five different regions of the sample shown in SSB images in Fig. 4b. Each region has a different vacancy configuration, labeled from zero to five, and a model for each is shown in Fig. 4a. The experimental ratios are shown in Fig. 4c alongside the values calculated from the DFT- and IAM-based simulations. Both experimental and DFT configurations are presented in their respective plots in order of increasing vacancy density around the measured . Of the experimental configurations, zero has the lowest average number of vacancies around the and five has the highest.
The low doses used in the experiment produce significant error bars particularly for the experimental configurations with very few such as configuration five with only five . However the overall effect seems clear: as the density of vacancies around the increases so too does the ratio , entirely consistent with the findings of our theoretical calculations. Experimental configuration five contains only double vacancies, and they are next to each other, meaning the are either next to one or two double vacancies. The calculated ratio for this configuration is the highest, consistent with those of the DFT configurations containing divacancies. Experimental configuration zero, on the other hand, contains no divacancies and has a much lower ratio, more consistent with the DFT calculations containing only monovacancies.
Clearly we are hindered in our ability to quantify the charge transfer itself in our experiments at the level of precision imposed by the dose required here to avoid damaging the defects, as well as the small sample sizes used. Furthermore the experimental structures are generally part of a larger, more complex structural environment than captured by our DFT calculations. Nonetheless, the increase in the mean experimental ratios with the density of vacancies at the measured sites fits very well with those predicted from theory. It therefore seems clear we can indeed detect the charge transfer caused by defects and indeed are sensitive to their density.
3 Conclusions